OpenStudy (anonymous):

Solve two-fifths plus one-ninth. Then use estimation to check for reasonableness. Show your work.

OpenStudy (nottim):

2/5+1/9

OpenStudy (nottim):

Common multiple ( i think) is 45.

OpenStudy (nottim):

you COULD get the denominators to equal each other.

OpenStudy (nottim):

OR you COULD add them as decimals (if your teacher permits)

OpenStudy (anonymous):

Pretty straightforward.\[\frac{2}{5}+\frac{1}{9}=\frac{18}{45}+\frac{5}{45}=\frac{23}{45}\] Two fifths is four tenths plus one ninth (a little more than a tenth) would give a little more than a half. 23/45 is a little more than a half.